# Towards Efficient Training and Evaluation of Robust Models against $l_{0}$ Bounded Adversarial Perturbations

### v1

@@ 4.2 Sparse-AutoAttack (sAA)
4.3 Adversarial Training
@@
In addition to robustness evaluation, we also explore adversarial training to build robust models against sparse perturbations. In the framework of adversarial training, the attack is used to generate adversarial perturbation in each training iteration, so the attack algorithm should not be too computationally expensive. In this regard, we run the untargeted sPGD (Algorithm 1) for $20$ iterations to generate sparse adversarial perturbations during training. We incorporate sPGD in the framework of vanilla adversarial training [4] and TRADES [20] and name corresponding methods sAT and sTRADES, respectively. Note that we use sAT and sTRADES as two examples of applying sPGD to adversarial training, since sPGD can be incorporated into any other adversarial training variant as well. To accommodate the scenario of adversarial training, we make the following modifications to sPGD.
@@
Multi-$k$ Strategy: Inspired by $l_{1}$-APGD [18] and Fast-EG-$l_{1}$ [19], multi-$k$ strategy is adopted to strengthen the robustness of model. That is, we use a larger sparsity threshold, i.e., $k$ in Algorithm 1, in the training phase than in the test phase.
@@ 5 Experiments
In this section, we conduct extensive experiments to compare our attack methods with baselines in evaluating the robustness of various models against $l_{0}$ bounded perturbations. Besides the effectiveness with an abundant query budget, we also study the efficiency of our methods when we use limited iterations to generate adversarial perturbations. Our results demonstrate that the proposed sPGD achieves the best performance among white-box attacks. When using limited iterations, sPGD achieves significantly better performance than existing methods. Therefore, sAA, consisting of the best white-box and black-box attacks, has the best attack success rate. sPGD, due to its efficiency, is utilized for adversarial training to obtain the best robust models against $l_{0}$ bounded adversarial perturbations. To further demonstrate the efficacy of our method, we evaluate the transferability of sPGD. The results show that sPGD has a high transfer success rate, making it applicable in more practical scenarios. In addition, we conduct ablation studies for analysis. The adversarial examples generated by our methods are visually interpretable and presented in Appendix C. Implementation details are deferred to Appendix B.
Table 1: Robust accuracy of various models on different attacks that generate $l_{0}$ bounded perturbations, where the sparsity level $k=20$. The models are trained on CIFAR-10. Note that we report results of Sparse-RS (RS) with fine-tuned hyperparameters, which outperforms its original version in Croce et al. [16]. CornerSearch (CS) is evaluated on $1000$ samples due to its high computational complexity.
Model Network Clean Black-Box White-Box sAA
CS RS SF PGD0 SAIF sPGDproj sPGDunproj
@@
5.1 Evaluation of Different Attack Methods
First, we compare our proposed sPGD, including sPGDproj and sPGDunproj as defined in Section 4.2, and sAA with existing white-box and black-box attacks that generate sparse perturbations. We evaluate different attack methods based on the models trained on CIFAR-10 [59] and report the robust accuracy with $k=20$ on the whole test set in Table 1. Additionally, the results on ImageNet100 [60] and a real-world traffic sign dataset GTSRB [61] are reported in Table 2. Note that the image sizes in GTSRB vary from $15\times 15$ to $250\times 250$. For convenience, we resize them to $224\times 224$ and use the same model architecture as in ImageNet-100. Furthermore, only the training set of GTSRB has annotations, we manually split the original training set into a test set containing $1000$ instances and a new training set containing the rest data. In Appendix A.1, we report more results on CIFAR-10 with $k=10$, $k=15$, and the results of models trained on CIFAR-100 [59] in Table 7, 8, 9, respectively to comprehensively demonstrate the effectiveness of our methods.
Table 2: Robust accuracy of various models on different attacks that generate $l_{0}$ bounded perturbations. ResNet34 (RN-34) [62] is used as the network architecture. Our sAT model is trained with the sparsity level $k=1200$ and the iteration number of the attack $t=20$. (a) The models are trained on ImageNet-100 [60], and the $k$ for attacks is set to $200$. (b) The models are trained on GTSRB [61], and the $k$ for attacks is set to $600$. Note that we report results of Sparse-RS (RS) with tuned hyperparameters, which outperforms its original version in Croce et al. [16]. All methods are evaluated on $500$ samples, and CornerSearch (CS) is not evaluated here due to its high computational complexity, i.e. nearly 1 week on one GPU for each run.
Model Network Clean Black-Box White-Box sAA
CS RS SF PGD0 SAIF sPGDproj sPGDunproj
Vanilla RN-34 83.0 - 0.2 5.8 9.8 0.6 0.2 0.4 0.0
$l_{1}$-adv. trained, $\epsilon=72$
Fast-EG-$l_{1}$ RN-34 69.2 - 50.2 43.4 50.2 43.0 18.6 19.0 16.6
$l_{0}$-adv. trained, $k=200$
PGD0-A RN-34 76.0 - 6.8 57.2 37.4 11.0 1.8 18.8 1.8
sAT RN-34 86.2 - 61.4 84.2 83.0 69.0 78.0 77.8 61.2
(a) ImageNet, $k=200$
Model Network Clean Black-Box White-Box sAA
CS RS SF PGD0 SAIF sPGDproj sPGDunproj
Vanilla RN-34 99.9 - 18.0 18.9 63.7 9.5 0.3 0.3 0.3
$l_{0}$-adv. trained, $k=600$
PGD0-A RN-34 99.8 - 37.6 23.6 59.8 13.8 0.0 0.0 0.0
sAT RN-34 99.8 - 88.4 96.8 99.8 96.2 88.6 96.2 85.4
(b) GTSRB, $k=600$
@@ 5.1 Evaluation of Different Attack Methods
Attacks: We compare our methods with various existing black-box and white-box attacks that generate $l_{0}$ bounded perturbations. The black-box attacks include CornerSearch (CS) [13] and Sparse-RS (RS) [16]. The white-box attacks include SparseFool (SF) [12], PGD0 [13] and Sparse Adversarial and Interpretable Attack Framework (SAIF) [36]. The implementation details of each attack are deferred to Appendix B. Specifically, to exploit the strength of these attacks in reasonable running time, we run all these attacks for either $10000$ iterations or the number of iterations where their performances converge. Note that, the number of iterations for all these attacks are no smaller than their default settings. In addition, we report the results of RS with fine-tuned hyperparameters, which outperforms its default settings in [16]. Finally, we report the robust accuracy under CS attack based on only $1000$ random test instances due to its prohibitively high computational complexity.
@@
(a) CIFAR-10, $k=20$
(b) ImageNet-100, $k=200$

### v3

@@ 3.2 Sparse-AutoAttack (sAA)
4 Structured Sparse Adversarial Attack
In this section, we extend Sparse-PGD (sPGD) and Sparse-AutoAttack (sAA) to generate structured sparse perturbations.
4.1 Formulation of Structured Sparsity
Given an input ${\bm{x}}\in\mathbb{R}^{d}$, we can partition its $d$ features into several groups that may overlap and define structured sparsity based on them. These groups can represent pixels of a row, a column, a patch, or a particular pattern. Without the loss of generality, we use $\{1,2,...,d\}$ as the indices of $d$ features of the input and $N$ set of indices ${\mathcal{G}}=\{G_{j}\}_{j=1}^{N}$ to represent the groups. For any perturbation, we define its group $l_{0}$ norm based on groups ${\mathcal{G}}$ as the minimal number of groups needed to cover its non-zero components. Mathematically, Bach [24] proposed a convex envelope of group $l_{0}$ norm as follows:
$\Omega({\bm{x}})=\sum_{k=1}^{d}|x_{\pi_{k}}|\left[F(\{\pi_{1},...,\pi_{k}\})-F(\{\pi_{1},...,\pi_{k-1}\})\right].$ (7)
where $F$ is a submodular function defined on the subsets of $V=\{1,2,...,d\}$ and $\{\pi_{i}\}_{i=1}^{d}$ is a permutation of $\{1,2,...,d\}$. More specifically, $F(A)$ indicates the minimal number of groups from ${\mathcal{G}}$ to cover the set $A$ and $\{\pi_{i}\}_{i=1}^{d}$ indicates the components of $|{\bm{x}}|$ in the decreasing order, i.e., $|x_{\pi_{1}}|\geq...\geq|x_{\pi_{d}}|\geq 0$. Compared with the convex envelope $\Omega$, the group $l_{0}$ norm $\Omega_{0}$ is defined as follows:
$\Omega_{0}({\bm{x}})$ $=\sum_{k=1}^{d}\mathbbm{1}(|x_{\pi_{k}}|)\left[F(\{\pi_{1},...,\pi_{k}\})-F(\{\pi_{1},...,\pi_{k-1}\})\right]$ (8)
$=F(\{\pi_{1},...,\pi_{d^{\prime}}\})\ \ \mathrm{where}\ x_{\pi_{d^{\prime}}}\neq 0\ \mathrm{and}\ x_{\pi_{d^{\prime}+1}}=0.$
Similar to the difference between $l_{1}$ norm and $l_{0}$ norm, we replace $|x_{\pi_{k}}|$ with $\mathbbm{1}(|x_{\pi_{k}}|)$ in Eq. (8) to indicate the number of groups with non-zero entities, where $\mathbbm{1}$ is an indicator function. The second equality in Eq. (8) is based on the decreasing order of $\{|x_{\pi_{i}}|\}_{i=1}^{d}$ and is a more straightforward definition of the group $l_{0}$ norm. Like the $l_{0}$ norm, the adversarial budgets based on the group $l_{0}$ norm are not convex, either.
In general, it is difficult to decide the minimum number of groups to cover non-zero elements, i.e., to calculate the function $F$. Therefore, it is challenging to directly apply sPGD to the group $l_{0}$ norm constraints. To address this issue, we can approximate $\Omega_{0}$ by a tight surrogate that facilitates the optimization by sPGD. In this regard, we introduce a binary group mask ${\bm{v}}=[v_{1},...,v_{N}]\in\{0,1\}^{N}$ to indicate whether a specific group is chosen to be perturbed. Mathematically, $\forall i$, if ${\bm{x}}_{i}\neq 0$, then $\exists j\in\{1,2,...,N\}$, $v_{j}=1$ and $i\in G_{j}$. In this context, we propose the following approximation of the group $l_{0}$ norm as the surrogate:
$\Omega_{0}^{\prime}({\bm{x}},{\bm{v}})=\sum_{i=1}^{N}\mathbbm{1}(\|{\bm{x}}_{G_{i}}\|_{p})\cdot v_{i},$ (9)
where ${\bm{x}}_{G_{i}}$ is the subvector of ${\bm{x}}$ on the indices in $G_{i}$, and $p\in[0,+\infty]$. Compared with $\Omega_{0}$ in Eq. (8), $\Omega^{\prime}_{0}$ in Eq. (9) does not require the number of groups in ${\mathcal{G}}$ to cover non-zero elements in ${\bm{x}}$ to be minimum. Instead, $\Omega^{\prime}_{0}$ calculates the number of perturbed groups in ${\mathcal{G}}$ selected by the binary vector ${\bm{v}}$. In addition, when $\mathbbm{1}(\|{\bm{x}}_{G_{i}}\|_{p})=1$ for all $i\in\{j~{}|~{}v_{j}=1\}$, the approximated group $l_{0}$ norm $\Omega_{0}^{\prime}({\bm{x}},{\bm{v}})$ is equivalent to the $l_{0}$ norm of the group mask ${\bm{v}}$, i.e., $\Omega_{0}^{\prime}({\bm{x}},{\bm{v}})=\|{\bm{v}}\|_{0}$. In summary, we have the following theorem:
Theorem 4.1.
Given an input ${\bm{x}}\in\mathbb{R}^{d}$, a set of groups $\mathcal{G}=\{G_{j}\}_{j=1}^{N}$ and any vector ${\bm{v}}$ satisfying the constraint in the definition of $\Omega^{\prime}_{0}$, then we have $\Omega_{0}({\bm{x}})\leq\Omega_{0}^{\prime}({\bm{x}},{\bm{v}})\leq\|{\bm{v}}\|_{0}$.
Proof.
The first inequality is based on the minimum optimality condition of the function $F$ in Eq. (8). Since $\Omega_{0}({\bm{x}})$ indicates the minimal number of groups needed to cover the indices of non-zero elements in ${\bm{x}}$, we have at least the same number of terms in Eq. (9) where ${\bm{x}}_{G_{i}}\neq 0$ and $v_{i}=1$. Therefore, we have $\Omega_{0}({\bm{x}})\leq\Omega^{\prime}_{0}({\bm{x}})$. For the second inequality, ${\bm{v}}$ is a binary vector, i.e., $\forall i$, ${\bm{v}}_{i}\in\{0,1\}$, so $\Omega_{0}^{\prime}({\bm{x}},{\bm{v}})\leq\|{\bm{v}}\|_{0}$ is clear.
Figure 1: The ratio between the group $l_{0}$ norm and the approximated group $l_{0}$ norm. The x-axis is the approximated group $l_{0}$ norm, and the y-axis is the ratio. We plot the ratios of random and adversarial $3\times 3$ patch perturbations, where the adversarial perturbations are generated on vanilla ResNet-18 trained on CIFAR-10. The solid line and shadow denote the mean value and standard deviation, respectively.
Theorem 4.1 indicates that the number of perturbed groups determined by ${\bm{v}}$ can be larger than the minimal number of groups to cover the perturbed features ${\bm{x}}$, i.e., $\Omega_{0}^{\prime}({\bm{x}},{\bm{v}})$ can be larger than $\Omega_{0}({\bm{x}})$. The gap between $\Omega_{0}^{\prime}({\bm{x}},{\bm{v}})$ and $\Omega_{0}({\bm{x}})$ stem from the potential overlap among groups in ${\mathcal{G}}$. However, we demonstrate that $\Omega_{0}^{\prime}({\bm{\delta}},{\bm{v}})$ is quite close to $\Omega_{0}({\bm{\delta}})$ in practice. To corroborate this, we calculate the ratio between the group $l_{0}$ norm and approximated group $l_{0}$ norm, i.e., $\Omega_{0}({\bm{\delta}})/\Omega_{0}^{\prime}({\bm{\delta}},{\bm{v}})$ of both random and adversarial perturbations under different sparsity level $\epsilon$ defined by $\Omega_{0}^{\prime}({\bm{\delta}},{\bm{v}})\leq\epsilon$. As illustrated in Figure 1, when applying $3\times 3$ patch perturbations, we can observe that the ratio $\Omega_{0}({\bm{\delta}})/\Omega_{0}^{\prime}({\bm{\delta}},{\bm{v}})$ decreases as $\epsilon$ increases. Specifically, when $\epsilon$ is smaller than $5$, the ratios of both random and adversarial perturbations are larger than $0.95$. Notably, the $l_{0}$ norm of $5$ non-overlapping $3\times 3$ patches reaches $45$, which is very large for a sparse perturbation used in practice, so we can conclude based on the simulation results that $\Omega_{0}^{\prime}$ is a good approximation of $\Omega_{0}$ in the context of the structured sparse attack.
4.2 sPGD for Structured Sparse Perturbations
Figure 2: Pipeline of sPGD for structured sparse perturbations. The continuous group mask $\widetilde{{\bm{v}}}$ is first projected to get the binary group mask ${\bm{v}}$ to ensure $\|{\bm{v}}\|_{0}\leq\epsilon$, which is similar to Eq. (6). Given the kernel ${\bm{k}}\in\{0,1\}^{r\times r}$ with a customized pattern, we can transform ${\bm{v}}$ to the pixel mask ${\bm{m}}$ using transposed convolution and clipping (see Eq. (10)). Finally, we element-wisely multiply ${\bm{m}}$ with the dense perturbation ${\bm{p}}$ to obtain the structured sparse perturbation ${\bm{\delta}}$. Note that the continuous group mask $\widetilde{{\bm{v}}}$ and the dense perturbation ${\bm{p}}$ are learnable.
Based on the approximated group $l_{0}$ norm $\Omega_{0}^{\prime}$ proposed in Eq. (9), we extend Sparse-PGD (sPGD) to generate structured sparse perturbations. Similarly, we decompose the perturbation ${\bm{\delta}}$ into a magnitude tensor ${\bm{p}}$ and a binary pixel mask ${\bm{m}}$. Like unstructured cases, the magnitude tensor ${\bm{p}}$ is updated using $l_{\infty}$ PGD, as in Eq. (4).
Based on the definition of structured sparsity and $\Omega^{\prime}_{0}$ defined in Eq. (9), if a specific group $G_{i}$ is chosen to be perturbed, then we let $v_{i}=1$, otherwise $v_{i}$ will be set 00. In this regard, finding the optimal ${\bm{m}}$ s.t. $\Omega^{\prime}_{0}({\bm{m}},{\bm{v}})\leq\epsilon$ is equivalent to finding the optimal group mask ${\bm{v}}$ s.t. $\|{\bm{v}}\|_{0}\leq\epsilon$, which can be resolved by sPGD. Similar to the unstructured cases, we optimize the continuous alternative $\widetilde{{\bm{v}}}$ rather than optimizing a binary ${\bm{v}}$ directly. In the following, we will elaborate on the mapping from ${\bm{v}}$ to ${\bm{m}}$ using different examples.
Column / Row: If we aim to perturb at most $\epsilon$ columns in an image ${\bm{x}}\in[0,1]^{h\times w\times c}$. Since we can partition ${\bm{x}}$ into $w$ groups (i.e., columns), ${\bm{v}}$ has the size of $1\times w$. Thus, ${\bm{m}}$ can obtained through expanding ${\bm{v}}$ from $1\times w$ to $h\times w$. As for row, the only difference is that the size of ${\bm{v}}$ is $h\times 1$. Note that, there is no overlap between any two columns or rows, so $\Omega^{\prime}_{0}({\bm{m}},{\bm{v}})\equiv\Omega_{0}({\bm{m}})$ in this case.
Patch / Any Pattern: Without loss of generality, we assume that the perturbations are contained to be localized in $r\times r$ patches. To avoid wasting budget due to the potential overlap at the border or corner of images, we let ${\bm{v}}\in\{0,1\}^{(h-r+1)\times(w-r+1)}$, and we can derive the mapping operation from ${\bm{v}}$ to ${\bm{m}}$ as follows:
${\bm{m}}=\min\left(\mathrm{TConv}({\bm{v}},{\bm{k}},s),1\right),$ (10)
where $\mathrm{TConv}$ is the transposed convolution operation, ${\bm{k}}=\mathbf{1}^{r\times r}$ is the kernel with all-one entities, $s=1$ denotes the stride. The clipping operation is to ensure the output value assigned to ${\bm{m}}$ is binary. Moreover, we can customize the kernel ${\bm{k}}\in\{0,1\}^{r\times r}$ in Eq. (10) to make the perturbations localized in groups with any desired patterns, such as hearts, stars, letters and so on. Notably, when ${\bm{k}}=[1]$, i.e., a $1\times 1$ matrix with a sole value $1$, the resulting perturbations are degraded to unstructured sparse ones; when ${\bm{k}}=\mathbf{1}^{h\times 1}$ or $\mathbf{1}^{1\times w}$, specific columns or rows are perturbed. The pipeline of sPGD for structured sparse perturbations in the general case is illustrated in Figure 2.
In addition to the forward mapping operation defined in Eq. (10), we provide its backward function as follows:
${\bm{g}}_{{\bm{v}}}=\mathrm{Conv}({\bm{g}}_{{\bm{m}}},{\bm{k}},s),$ (11)
where $\mathrm{Conv}$ is the convolution operation, and ${\bm{g}}_{{\bm{m}}}$ is the gradient of ${\bm{m}}$. We neglect the clipping operation when calculating the gradient of ${\bm{v}}$ since it is not always differentiable.
The extension of sPGD for structured sparse perturbations offers a unified framework to generate sparse adversarial perturbations. We provide its pseudo-code in Algorithm 2.
Algorithm 2 Sparse-PGD for Structured Sparse Perturbations
1:  Input: Clean image: ${\bm{x}}\in[0,1]^{h\times w\times c}$; Model parameters: ${\bm{\theta}}$; Customized pattern kernel: ${\bm{k}}\in\{0,1\}^{r\times r}$; Stride: $s=1$; Max iteration number: $T$; Tolerance: $t$; group $l_{0}$ budget: $\epsilon$; Step size: $\alpha$, $\beta$; Small constant: $\gamma=2\times 10^{-8}$
2:  Random initialize ${\bm{p}}$ and $\widetilde{{\bm{v}}}$
3:  ${\bm{v}}=\Pi_{{\mathcal{S}}_{{\bm{v}}}}(\sigma(\widetilde{{\bm{v}}}))$ where ${\mathcal{S}}_{{\bm{v}}}=\{{\bm{v}}|\|{\bm{v}}\|_{0}\leq\epsilon\}$.
4:  ${\bm{m}}=\min\left(\mathrm{TConv}({\bm{v}},{\bm{k}},s),1\right)$
5:  for $i=0,1,...,T-1$ do
6:     Calculate the loss $\mathcal{L}({\bm{\theta}},{\bm{x}}+{\bm{p}}\odot{\bm{m}})$
7:     if unprojected then
8:        $g_{{\bm{p}}}=\nabla_{{\bm{\delta}}}{\mathcal{L}}\odot\min\left(\mathrm{TConv}(\sigma(\widetilde{{\bm{v}}}),{\bm{k}},s),1\right)$
9:     else
10:        $g_{{\bm{p}}}=\nabla_{\bm{\delta}}{\mathcal{L}}\odot{\bm{m}}$
11:     end if
12:     $g_{\widetilde{{\bm{v}}}}=\mathrm{Conv}(\nabla_{{\bm{m}}}{\mathcal{L}},{\bm{k}},s)\odot\sigma^{\prime}(\widetilde{{\bm{v}}})$
13:     ${\bm{p}}=\Pi_{{\mathcal{S}}_{{\bm{p}}}}({\bm{p}}+\alpha\cdot\mathtt{sign}(g_{{\bm{p}}}))$
14:     ${\bm{d}}=g_{\widetilde{{\bm{v}}}}/\|g_{\widetilde{{\bm{v}}}}\|_{2}$ if $||g_{\widetilde{{\bm{v}}}}||_{2}\geq\gamma$ else 00
15:     ${\bm{v}}_{old},\ \widetilde{{\bm{v}}}={\bm{v}},\ \widetilde{{\bm{v}}}+\beta\cdot{\bm{d}}$
16:     ${\bm{v}}=\Pi_{{\mathcal{S}}_{{\bm{v}}}}(\sigma(\widetilde{{\bm{v}}}))$
17:     ${\bm{m}}=\min\left(\mathrm{TConv}({\bm{v}},{\bm{k}},s),1\right)$
18:     if attack succeeds then
19:        break
20:     end if
21:     if $||{\bm{v}}-{\bm{v}}_{old}||_{0}\leq 0$ for $t$ consecutive iters then
22:        Random initialize $\widetilde{{\bm{v}}}$
23:     end if
24:  end for
25:  Output: Perturbation: ${\bm{\delta}}={\bm{p}}\odot{\bm{m}}$
4.3 sAA for Structured Sparse Perturbations
Similar to Section 3.2, we also combine sPGDunproj, sPGDproj and Sparse-RS to build Sparse-AutoAttack (sAA) for reliable evaluation of robustness against structured sparse perturbations. We extended the original version of Sparse-RS so that it can generate structured sparse perturbations with any customized patterns. Specifically, we just apply a pattern mask to the patches generated by Sparse-RS.
5 Adversarial Training
@@
In addition to adversarial attacks, we explore adversarial training to build robust models against sparse perturbations. In the framework of adversarial training, the attack is used to generate adversarial perturbation in each training iteration, so the attack algorithm should not be too computationally expensive. In this regard, we run the untargeted sPGD (Algorithm 1) for $20$ iterations to generate sparse adversarial perturbations during training. We incorporate sPGD in the framework of vanilla adversarial training [4] and TRADES [25] and name corresponding methods sAT and sTRADES, respectively. Note that we use sAT and sTRADES as two examples of applying sPGD to adversarial training since sPGD can be incorporated into any other adversarial training variant as well. To accommodate the scenario of adversarial training, we make the following modifications to sPGD.
@@
Multi-$\epsilon$ Strategy: Inspired by $l_{1}$-APGD [19] and Fast-EG-$l_{1}$ [20], multi-$\epsilon$ strategy is adopted to strengthen the robustness of model. That is, we use a larger sparsity threshold, i.e., $\epsilon$ in Algorithm 1, in the training phase than in the test phase.
@@ 6 Experiments
In this section, we conduct extensive experiments to compare our attack methods with baselines in evaluating the robustness of various models against $l_{0}$ bounded and structured sparse perturbations. Besides the effectiveness with an abundant query budget, we also study the efficiency of our methods. Our results demonstrate that our sPGD performs best among white-box attacks. With limited iterations, sPGD achieves significantly better performance than existing methods. Therefore, sAA, consisting of the best white-box and black-box attacks, has the best attack success rate. sPGD, due to its efficiency, is utilized for adversarial training to obtain the best robust models against $l_{0}$ bounded and structured sparse adversarial perturbations. To further demonstrate the efficacy of our method, we evaluate the transferability of sPGD. The results show that sPGD has a high transfer success rate, making it applicable in practical scenarios. The adversarial examples generated by our methods are presented in Sec. 6.7. In addition, we conduct ablation studies for analysis in C. Implementation details are deferred to Appendix A.
6.1 Robustness against Unstructured Sparse Perturbations
Table 1: Robust accuracy of various models on different sparse attacks, where the sparsity level $\epsilon=20$. The models are trained on CIFAR-10. Note that we report results of Sparse-RS (RS) with fine-tuned hyperparameters, which outperforms its original version in Croce et al. [17]. CornerSearch (CS) is evaluated on $1000$ samples due to its high computational complexity.
Model Network Clean Black White sAA
CS RS SF PGD0 SAIF sPGDp sPGDu
@@ 6.1 Robustness against Unstructured Sparse Perturbations
First, we compare our proposed sPGD, including sPGDproj (sPGDp) and sPGDunproj (sPGDu) as defined in Section 3.2, and sAA with existing white-box and black-box attacks that generate $l_{0}$ bounded sparse perturbations. We evaluate different attack methods based on the models trained on CIFAR-10 [64] and report the robust accuracy with $\epsilon=20$ on the whole test set in Table 1. Additionally, the results on ImageNet100 [65] and a real-world traffic sign dataset GTSRB [66] are reported in Table 2. Note that the image sizes in GTSRB vary from $15\times 15$ to $250\times 250$. For convenience, we resize them to $224\times 224$ and use the same model architecture as in ImageNet-100. Furthermore, only the training set of GTSRB has annotations, we manually split the original training set into a test set containing $1000$ instances and a new training set containing the rest data. In Appendix B, we report more results on CIFAR-10 with $\epsilon=10$, $\epsilon=15$, and the results of models trained on CIFAR-100 [64] in Table 7, 8, 9, respectively, to demonstrate the efficacy of our methods.
@@
Attacks: We compare our methods with various existing black-box and white-box attacks that generate $l_{0}$ bounded perturbations. The black-box attacks include CornerSearch (CS) [13] and Sparse-RS (RS) [17]. The white-box attacks include SparseFool (SF) [12], PGD0 [13] and Sparse Adversarial and Interpretable Attack Framework (SAIF) [41]. The implementation details of each attack are deferred to Appendix A. Specifically, to exploit the strength of these attacks in reasonable running time, we run all these attacks for either $10000$ iterations or the number of iterations where their performances converge. Note that, the number of iterations for all these attacks are no smaller than their default settings. In addition, we report the results of RS with fine-tuned hyperparameters, which outperforms its default settings in [17]. Finally, we report the robust accuracy under CS attack based on only $1000$ random test instances due to its prohibitively high computational complexity.
@@
Table 2: Robust accuracy of various models on different sparse attacks. Our sAT model is trained with $\epsilon=1200$. (a) Results on ImageNet-100 [65]. (b) Results on GTSRB [66]. Note that the results of Sparse-RS (RS) with tuned hyperparameters are reported. All models are RN-34, and are evaluated on $500$ samples. The results of SparseFool (SF) and PGD0 are included due to their poor performance, and CornerSearch (CS) is not evaluated here due to its high computational complexity, i.e. nearly 1 week on one GPU for each run.
Model Clean Black White sAA
RS SAIF sPGDp sPGDu
Vanilla 83.0 0.2 0.6 0.2 0.4 0.0
$l_{0}$-adv. trained, $\epsilon=200$
PGD0 76.0 6.8 11.0 1.8 18.8 1.8
sAT 86.2 61.4 69.0 78.0 77.8 61.2
(a) ImageNet-100, $\epsilon=200$
Model Clean Black White sAA
RS SAIF sPGDp sPGDu
Vanilla 99.9 18.0 9.5 0.3 0.3 0.3
$l_{0}$-adv. trained, $\epsilon=600$
PGD0 99.8 37.6 13.8 0.0 0.0 0.0
sAT 99.8 88.4 96.2 88.6 96.2 85.4
(b) GTSRB, $\epsilon=600$
@@
6.2 Robustness against Structured Sparse Perturbations
Table 3: Robust accuracy of various models on different structured sparse attacks. (a) Results on CIFAR-10, all models are PRN-18. (b) Results on ImageNet-100, the test set contains $500$ samples, all models are RN-34. Note that the evaluated group sparsity levels are approximately equivalent to the $l_{0}$ sparsity levels used in Table 1 and 2, and the Sparse-RS (RS) used here is the proposed extended version.
Model Clean Black White sAA
RS LOAP sPGDp sPGDu
row, $\epsilon=1$
Vanilla 93.9 22.5 - 1.0 1.3 1.0
sTRADES-$l_{0}$ 89.8 83.1 - 54.6 65.0 53.9
sTRADES-row 89.3 85.9 - 77.3 83.7 77.3
$3\times 3$ patch, $\epsilon=2$
Vanilla 93.9 6.7 4.5 1.9 6.6 1.7
sTRADES-$l_{0}$ 89.8 65.5 60.0 46.9 67.5 46.7
sTRADES-p3x3 87.9 77.5 81.4 72.8 81.7 72.3
$5\times 5$ patch, $\epsilon=1$
Vanilla 93.9 7.5 2.3 1.7 3.2 1.7
sTRADES-$l_{0}$ 89.8 56.4 35.0 26.7 44.3 26.7
sTRADES-p5x5 89.5 72.9 73.1 56.0 73.3 55.9
(a) CIFAR-10
Model Clean Black White-Box sAA
RS LOAP sPGDp sPGDu
row, $\epsilon=1$
Vanilla 83.0 54.8 - 5.2 8.0 5.2
sAT-$l_{0}$ 86.2 78.6 - 65.8 73.8 65.8
sAT-row 81.0 78.2 - 76.6 78.4 76.6
$10\times 10$ patch, $\epsilon=2$
Vanilla 83.0 13.0 4.6 4.6 6.4 4.0
sAT-$l_{0}$ 86.2 27.8 10.8 10.6 33.0 10.0
sAT-p10x10 81.6 64.0 62.8 35.8 73.2 35.8
$14\times 14$ patch, $\epsilon=1$
Vanilla 83.0 7.5 2.3 1.7 12.0 1.7
sAT-$l_{0}$ 86.2 28.3 10.2 15.4 34.8 14.6
sAT-p14x14 80.0 57.2 73.0 41.2 71.2 39.6
(b) ImageNet-100
Apart from unstructured sparse perturbations, we evaluate the effectiveness of our method in Algorithm 2 in generating structured sparse perturbations in this subsection. Given the paucity of available methods for comparison, we focus on comparing our method with the extended Sparse-RS (RS) and LOAP [16], which is a white-box approach generating adversarial patches, on CIFAR-10 and ImageNet-100. For a comprehensive evaluation, we include the results of different types of structured sparse perturbations, e.g., row and patches with different group sparsity level and different sizes. It should be noted that the unstructured $l_{0}$ norms of the structured perturbations studied here are approximately the same as the perturbations in Table 1 and 2. In addition, we evaluate the attacks on vanilla models, models trained with unstructured $l_{0}$ bounded perturbations, and those trained with the specific structured sparse perturbations.
The results in Table 3 indicate that the proposed sAA also achieves the state-of-the-art performance in generating structured sparse perturbations in almost all cases. Furthermore, the models trained with our methods exhibit the strongest robustness against structured sparse perturbations. As anticipated, models trained on specific structured sparse adversarial examples significantly outperform those trained on unstructured sparse adversarial examples. This highlights the inherent limitation of adversarially trained models in maintaining robustness against unseen types of perturbations during training.
6.3 Adversarial Watermarks
Table 4: Robust accuracy of vanilla model (RN-34) against adversarial watermarks with different $l_{\infty}$ adversarial budgets. The experiment is conducted on ImageNet-100.
$\epsilon_{\infty}$ $8/255$ $16/255$ $32/255$ $64/255$
RS 59.0 44.6 32.4 26.4
sPGDp 47.2 25.4 16.6 9.0
sPGDu 56.4 39.6 25.8 16.8
sAA 44.2 24.0 15.6 8.8
(a) $18\times 18$ circle
$\epsilon_{\infty}$ $8/255$ $16/255$ $32/255$ $64/255$
RS 14.2 12.8 9.2 5.8
sPGDp 3.8 2.6 2.4 2.2
sPGDu 7.4 4.8 2.8 2.0
sAA 2.0 0.8 0.8 0.6
(b) $60\times 60$ letter “A”
Although only a few pixels are perturbed in sparse attacks, such perturbations are still perceptible due to their unconstrained magnitude. In this subsection, we additionally constrain the $l_{\infty}$ norm of magnitude ${\bm{p}}$, meaning that the feasible set for $vp$ is rewritten as ${\mathcal{S}}_{\bm{p}}=\{{\bm{p}}\in\mathbb{R}^{h\times w\times c}|\|{\bm{p}}\|_{\infty}\leq\epsilon_{\infty},~{}0\leq{\bm{x}}+{\bm{p}}\leq 1\}$. Combining with structured sparsity constraint, we can generate alleged adversarial watermarks with our method.
To evaluate the performance of our method in generating adversarial watermarks, we conduct experiments with different patterns and different $\l_{\infty}$ adversarial budgets. From Table 4, we can find that when the size of the pattern is large, like $60\times 60$ letter “A” with $1747$ non-zero elements, our attack can still achieve decent performance under different $l_{\infty}$ adversarial budgets. In contrast, when we adopt a $18\times 18$ circle pattern with $208$ non-zero elements, the attack success rate abruptly declines as the $l_{\infty}$ adversarial budget decreases. Nevertheless, from the aspect of attack success rate, our approaches still outperform Sparse-RS by a large margin.
6.4 Comparison under Different Iteration Numbers and Different Sparsity Levels
(a) CIFAR-10, $\epsilon=20$
(b) ImageNet-100, $\epsilon=200$
(c) CIFAR-10, $t=10000$